OpenStudy (anonymous):

How do you integrate e^2x sin(3x)dx?

OpenStudy (bahrom7893):

Use integration by parts. Btw when is this due?

OpenStudy (anonymous):

its a revision, we cant figure it out

OpenStudy (bahrom7893):

No i just meant i have to do my own hw too, so do u need this like right away or by tonite or by 2morro mornin?

OpenStudy (bahrom7893):

ill just do it now

OpenStudy (anonymous):

we would just like to know what to do with the sin(3x)

OpenStudy (anonymous):

thanks!

OpenStudy (bahrom7893):

ima write this out on paper and email it to u. what's ur email?

OpenStudy (anonymous):

moca341@hotmail.com .. thank you :)

OpenStudy (bahrom7893):

np

OpenStudy (bahrom7893):

almost done...

OpenStudy (anonymous):

great!

OpenStudy (bahrom7893):

i will take separate pix and email them one by one, the whole solution won't fit lol

OpenStudy (anonymous):

thats fine! haha

OpenStudy (bahrom7893):

okay emailed it

OpenStudy (bahrom7893):

ask me if u have any questions, either here or by email and Fan if I helped!

OpenStudy (bahrom7893):

did u get it?

OpenStudy (anonymous):

It helped a lot, but i dont really get it!, ill try to figure it out! thanks a lot!

OpenStudy (bahrom7893):

no, ask me which part don't you understand?

OpenStudy (bahrom7893):

I think it might the place where I let that one integral be capital i

OpenStudy (anonymous):

i dont even understand the first line, dont worry about it!

OpenStudy (bahrom7893):

Oh okay so first I used integration by parts. See the stuff in the circle: I said Let u = e^(2x), then what is du/dx? du/dx = derivative of u = 2e^(2x), multiply both sides by dx to get: du = 2e^(2x)dx

OpenStudy (anonymous):

i dont see where the 1/3 cos (3x) comes from

OpenStudy (bahrom7893):

Oh okay that's just the integral of Sin(3x)

OpenStudy (bahrom7893):

I will work it out here: After that I said let dv/dx = Sin(3x), well then what's V? v is the integral of dv/dx with respect to x. V = Integral(dv/dx, dx) \[v = \int\limits_{}^{}Sin(3x)dx\]

OpenStudy (bahrom7893):

Here you have to use a simple substitution: Let a = 3x; then da = 3 dx. Now you have 3x in your integral, but you still need a 3 in front of dx. To do so, multiply and divide by 3, same as multiplying by 1

OpenStudy (anonymous):

ok, i get it, thanks a lot!

OpenStudy (bahrom7893):

\[v = \int\limits_{}^{}Sin(3x)*(3/3)*dx = (1/3) *\int\limits_{}^{}Sin(3x)*3*dx\]

OpenStudy (bahrom7893):

Now you have both a and da; your integral simplifies to: \[(1/3)\int\limits_{}^{}Sin(a)*da\]

OpenStudy (bahrom7893):

that's it then integrate as u would a regular sin and in the end replace a by 3x